OpenStudy (anonymous):

a=find the first, fourth, and eighth terms of each sequence. A(n)=-5*3^n-1

OpenStudy (anonymous):

\[A(1)=-5\times 3^{1-1}=-5\times 3^0=-5\] for the first one others are similar

OpenStudy (anonymous):

@satellite73 wait how did the 1 change to 0?

OpenStudy (anonymous):

the exponent is \(n-1\) if you replace \(n\) by \(1\) you get \(1-1=0\)